The length of a rectangle is 4 inches longer than its width. If the perimeter of the rectangle is 76​ inches, find the length
Answer by fractalier(6550)

Perimeter P = 2L + 2W = 76 and
L = W + 4
Now substitute and get
2(W+4) + 2W = 76
2W + 8 + 2W = 76
4W + 8 = 76
4W = 68
W = 17 inches
L = W + 4 = 21 inches